$1000 is deposited in a savings account that pays an annual rate of 6%, compounded monthly. Determine the value of the account in 2 1/2 years

Answer by stanbon(75887)

$1000 is deposited in a savings account that pays an annual rate of 6%, compounded monthly. Determine the value of the account in 2 1/2 years
---------
A(t) = P(1+(r/n))^(nt)
-------
A(2 1/2) = 1000(1 + (0.06/12))^(12(5/2))
----
A(2 1/2) = 1000(1.05)^30 = $1161.40
==============================
Cheers,
Stan H.